OpenStudy (anonymous):

Braden improves his score with each practice quiz he takes. The number of incorrect answers can be represented by the expression 20 – 5n where n represents the number of practice quizzes he has taken. How many answers will he get incorrect after taking 2 practice quizzes? 15 13 10 5

OpenStudy (anonymous):

Plug 2 in for n and then solve.

OpenStudy (anonymous):

so its 20-5x2

OpenStudy (anonymous):

Yep. Now just simplify that expression.

OpenStudy (anonymous):

15

hero (hero):

Let n be the number of quizzes. If n = 2, then 20 - 5n = 20 - 5(2) = 20 - 10 = 10 Therefore, if Brandon takes two quizzes, he will accumulate 10 incorrect answers.

OpenStudy (anonymous):

how its ten doe if half of 30 is 15

hero (hero):

What kind of math are you doing? You have to apply order of operations. There's a minus sign between the 20 and the 5n

hero (hero):

20 - 5n = 20 - 5(2) = 20 - 10 = 10

OpenStudy (anonymous):

i know that but that 5n yhue got to muiltiply the n which is 2 time 15 is 30 half is 15

hero (hero):

\[5 \times 2 = 10\]\[5 \times 2 \ne 15\]

OpenStudy (anonymous):

look 20-5=15x2=30 half =15

hero (hero):

I'm sorry, but you will need a math tutor to help you.

OpenStudy (anonymous):

yhue need a math tutor bihh

hero (hero):

If I needed a math tutor, I would be the one asking questions.

OpenStudy (anonymous):

look hoe get off mhy shidd

hero (hero):

I suppose you're looking to get abuse reported.

OpenStudy (anonymous):

ummhm